### v1

@@ 3.4.1 $\tau$-quantiles and the cross-section of stock returns
Table 7: Quantiles and the cross-section of stock returns in the U.S. This table shows the average monthly returns and annualised Sharpe ratios for decile long, short and long-short equal-weighted portfolios over the period 1995 to 2018. For each $\tau$-quantile forecast, the long portfolio is formed by taking the top 10% of stocks with the highest forecast each month, The short portfolio is formed by taking the bottom 10% of stocks with the lowest forecast each month. The long-short portfolio is created by taking the difference between the long and short portfolios. Values in brackets are t-statistics adjusted for Newey-West standard errors with 12 lags. Results are obtained separately for the full and liquid samples for the US only.
Full Sample Liquid Sample
Long Short Long-short Long Short Long-short
$\tau$ Mean SR Mean SR Mean SR Mean SR Mean SR Mean SR
0.01 1.22 (5.62) 1.67 0.14 (0.19) 0.04 1.09 (1.52) 0.31 0.95 (5.26) 1.10 0.16 (0.21) 0.05 0.79 (1.06) 0.24
0.05 1.25 (5.66) 1.76 -0.09 (-0.12) -0.02 1.34 (1.86) 0.39 0.96 (5.63) 1.15 0.15 (0.20) 0.04 0.81 (1.11) 0.25
0.1 1.28 (5.73) 1.79 -0.26 (-0.35) -0.07 1.53 (2.13) 0.45 0.98 (5.72) 1.17 0.11 (0.14) 0.03 0.87 (1.20) 0.26
0.2 1.34 (6.04) 1.87 -0.57 (-0.78) -0.16 1.91 (2.68) 0.57 1.02 (5.89) 1.21 0.04 (0.06) 0.01 0.98 (1.34) 0.30
0.3 1.46 (6.54) 1.96 -0.91 (-1.27) -0.27 2.37 (3.36) 0.74 1.10 (6.09) 1.25 -0.04 (-0.06) -0.01 1.14 (1.58) 0.36
0.4 1.88 (7.44) 2.06 -1.34 (-1.91) -0.42 3.22 (4.89) 1.12 1.39 (6.31) 1.34 -0.14 (-0.20) -0.04 1.53 (2.23) 0.54
0.5 2.88 (6.97) 1.78 -1.82 (-2.68) -0.62 4.70 (8.25) 2.27 1.86 (5.61) 1.10 -0.35 (-0.53) -0.12 2.21 (4.21) 1.21
0.6 3.44 (5.68) 1.37 -2.12 (-3.51) -0.88 5.56 (9.41) 3.96 1.74 (3.69) 0.71 -0.38 (-0.88) -0.21 2.11 (5.40) 1.34
0.7 3.48 (4.73) 1.08 -1.42 (-3.57) -1.10 4.90 (6.57) 1.99 1.66 (2.86) 0.56 0.48 (2.50) 0.54 1.18 (2.23) 0.43
0.8 2.98 (3.72) 0.82 0.44 (1.65) 0.62 2.55 (3.16) 0.75 1.28 (1.94) 0.39 0.71 (3.95) 0.85 0.57 (0.91) 0.18
0.9 2.12 (2.57) 0.55 0.75 (3.29) 1.10 1.37 (1.69) 0.38 0.91 (1.24) 0.27 0.76 (4.29) 0.93 0.15 (0.21) 0.05
0.95 1.71 (2.11) 0.45 0.86 (4.04) 1.22 0.85 (1.07) 0.23 0.77 (1.03) 0.23 0.80 (4.62) 0.98 -0.03 (-0.04) -0.01
0.99 1.36 (1.70) 0.36 0.97 (5.04) 1.22 0.39 (0.50) 0.11 0.60 (0.80) 0.17 0.84 (4.72) 1.01 -0.23 (-0.32) -0.07
@@
We show the highest average monthly returns and Sharpe ratios for the long-short portfolios formed based on the central $\tau$-quantiles, $\tau$s of $50$% or $60$%, for both the full sample and the liquid sample. Average monthly returns decrease as we move towards the tails of the distribution. For the full sample, average monthly returns are not significantly different from zero for $\tau$s greater than 80% and less than 10%. For the liquid sample, the average monthly returns fall even more sharply as one moves towards the tails of the distribution, becoming not significantly different from zero for $\tau$s greater than 70% and less than 40%. The highest Sharpe ratios are achieved by the long-short portfolios with the highest average monthly returns.
@@ 3.4.2 Distributional moments and the cross-section of stock returns
Table 8: Mean and median portfolios. This table shows average monthly returns and annualised Sharpe ratios for decile long-short equal-weighted portfolios based on the median and mean predictions from the two-stage quantile NN model, the mean prediction from the two-hidden-layer neural network (2hNN - Mean) and the median prediction from the two-stage NN model (2hNN - Median). The results are shown for the full and liquid samples and for individual regions, covering the period from 1995 to 2018. The long portfolio is formed by taking the top 10% of stocks with the highest forecast each month, and the short portfolio is formed by taking the bottom 10% of stocks with the lowest forecast each month. Values in parentheses are t-statistics adjusted for Newey-West standard errors with 12 lags. The significance of the difference in mean returns is tested using t-statistics with Newey-West standard errors with 12 lags.
@@
To benchmark our forecasts, we use the mean forecasts obtained directly from the feed-forward neural network model estimated with the mean squared error loss function. The long-short portfolio based on the mean forecast of our two-stage quantile model leads to significantly higher average monthly returns and Sharpe ratios compared to the portfolios based on the mean forecasts obtained directly from the simple neural network for the full sample across all regions. This is consistent with the $R^{2}$ results in Table 5, as our two-stage model outperforms the direct mean forecast in terms of $R^{2}$ across all regions and also leads to higher corresponding profitability. In the case of the full sample, the average monthly returns of our mean forecast are economically higher than the direct mean forecast in all regions and significantly higher for Japan, Asia-Pacific and globally.
Next, the average monthly returns of the mean- and median-based long-short portfolios are statistically indistinguishable for the liquid sample. This is also consistent with the $R^{2}$ results in Table 5, as there are no significant differences in the quality of the mean and median forecasts for the liquid sample. However, the results are different for the full sample. The difference between the mean and the median is associated with distributional asymmetry, which is more pronounced for stocks in the full sample. Mean-based long-short portfolios have significantly higher average monthly returns than median-based long-short portfolios. This is true for all regions and also globally. The difference between mean and median pricing for the full sample indicates the effect of distributional asymmetry in the cross-section of stock returns.
@@
The results on volatility show a negative relationship between predicted volatility and future return for the liquid sample, in line with previous academic research. However, the relationship is not statistically significant, which is to be expected given the focus on large cap stocks. The full sample results are plagued by market microstructure distortions for equal-weighted portfolios, such as bid-ask jumps, and the pattern is clearer for value-weighted portfolios. However, the effect remains statistically insignificant, which can be explained by our sample period, which focuses primarily on years more recent than the sample period of Ang et al. (2006), the original study that identified volatility as priced into the cross-section of returns.
@@ 4 Conclusion
We develop a novel machine learning approach to forecasting the full distribution of stock returns, focusing on both US and international markets. Our method uses a multi-head quantile neural network to predict a set of $\tau$ quantiles based on stock characteristics and market variables. We interpolate these quantiles using cubic B-splines to derive the full distribution and its moments, such as mean, volatility, skewness and kurtosis. This approach outperforms traditional parametric models and simpler neural networks in forecasting stock returns, providing more accurate out-of-sample mean and variance forecasts.
@@ B.1 Feed-forward neural networks
We use multi-head feed-forward neural networks with varying architectures as benchmarks for the two-stage model quantile forecasts. We consider a simple multi-output linear model without any hidden layers, neural networks with 1 hidden layer, and neural networks with 2 hidden layers as benchmarks for the two-stage model. Each model has 176 features and 37 outputs. The input features are the same as those for the standardized $\tau$-quantiles network in the two-stage model. The 37 outputs are the $\tau$-quantile forecasts of raw future stock returns for the same $\tau$ values as in the two-stage model. See Section 3.1.1 for the feature specification and Section 2 for the main model introduction.
Each neural network is trained using the aggregated multi-$\tau$-quantile loss.
@@
(b) One hidden layer neural network
@@
(c) Two hidden layers neural network
@@
(d) Single output neural network with two hidden layers
@@
1(a) depicts the linear model, 1(b) illustrates the neural network with one hidden layer, 1(c) shows the neural network with two hidden layers, and 1(d) presents the two hidden layers neural network with a single output node, which is used to generate mean forecasts directly using the mean squared error loss function. All networks have 176 input features; multi-output networks have 37 outputs, varying in the number of hidden layers and neurons.
@@ B.2 GARCH benchmark
We consider several variants of GARCH(1,1) model with the standardized residuals distributed according to normal and t-distributions. Mean $\mu$ is either estimated or set equal to the risk-free rate at the end of the estimation window plus 5% divided by 252. Assuming these distributions, we simulate 100,000 paths of residuals for 22 days ahead and consequently derive the corresponding return paths. During this process, volatility is simulated using Equation 7
@@
For the t-distribution, besides estimating the number of degrees of freedom from the data, we also consider fixing the number of degrees of freedom to 3, 4, and 5. Furthermore, we try GARCH(2, 2) and EGARCH(1, 1, 1) specifications also with t-distributed standardized residuals, in both cases estimating degrees of freedom from the data.
@@ Appendix C Additional empirical results
Figure C.1: The cross-sectional average of volatilities as a scaling factor. This figure shows the evolution over time of the cross-sectional average of stock-level volatilities for the full sample of U.S. stocks. Stock-level volatilities are calculated as exponential moving averages of squared daily returns with a smoothing factor of $0.94$. The scaling factor is used to standardize stock returns to be used as a label for the standardized $\tau$-quantile neural network or stage 1 of the two-stage model defined in Section 2.
@@
Median
Full Sample Liquid Sample
Decile USA Europe Japan Asia Pacific USA Europe Japan Asia Pacific
1 -1.84 (-2.11) -2.55 (-3.25) -1.46 (-1.89) -2.44 (-2.75) -0.53 (-0.75) -0.42 (-0.81) -0.87 (-1.45) -1.00 (-1.39)
2 -0.61 (-0.83) -1.07 (-1.75) -0.87 (-1.45) -1.42 (-2.04) 0.11 (0.25) 0.26 (0.61) -0.26 (-0.53) -0.01 (-0.02)
3 -0.21 (-0.34) -0.28 (-0.51) -0.27 (-0.49) -0.78 (-1.17) 0.48 (1.37) 0.48 (1.19) -0.12 (-0.25) 0.36 (0.68)
4 0.18 (0.34) 0.09 (0.24) -0.24 (-0.52) -0.83 (-1.27) 0.60 (2.11) 0.59 (1.69) -0.01 (-0.02) 0.44 (0.98)
5 0.50 (1.40) 0.44 (1.22) 0.03 (0.06) -0.34 (-0.55) 0.70 (2.46) 0.54 (1.56) 0.29 (0.70) 0.40 (1.00)
6 0.51 (1.52) 0.40 (1.07) -0.03 (-0.08) -0.08 (-0.16) 0.91 (3.73) 0.69 (2.16) 0.12 (0.36) 1.10 (2.75)
7 0.84 (3.00) 0.77 (2.31) 0.26 (0.64) 0.18 (0.38) 1.06 (4.01) 0.90 (2.62) 0.41 (1.09) 1.04 (3.00)
8 0.79 (2.58) 0.77 (2.51) 0.41 (1.03) 0.92 (1.94) 1.02 (3.39) 1.11 (3.24) 0.48 (1.46) 1.39 (3.71)
9 1.21 (4.17) 1.15 (3.01) 0.67 (1.66) 1.26 (2.94) 1.27 (4.59) 1.21 (3.27) 0.91 (2.52) 1.43 (3.76)
10 1.76 (5.06) 1.58 (3.90) 1.05 (2.49) 1.82 (3.87) 1.86 (4.94) 1.43 (3.58) 1.08 (2.69) 1.85 (4.31)
10-1 3.59 (5.02) 4.13 (6.86) 2.51 (4.86) 4.26 (7.67) 2.39 (4.43) 1.85 (4.42) 1.95 (5.10) 2.86 (5.70)

### v2

@@ 3.4.1 $\tau$-quantiles and the cross-section of stock returns
Table 7: Quantiles and the cross-section of stock returns in the U.S. This table shows the average and volatility of monthly returns and annualised Sharpe ratios for decile long, short and long-short equal-weighted portfolios over the period 1995 to 2018. For each $\tau$-quantile forecast, the long portfolio is formed by taking the top 10% of stocks with the highest forecast each month, The short portfolio is formed by taking the bottom 10% of stocks with the lowest forecast each month. The long-short portfolio is created by taking the difference between the long and short portfolios. Values in brackets are t-statistics adjusted for Newey-West standard errors with 12 lags. Results are obtained separately for the full and liquid samples for the US only.
Full Sample
Long Short Long-short
$\tau$ Mean Vol SR Mean Vol SR Mean Vol SR
0.01 1.22 (5.62) 2.53 1.67 0.14 (0.19) 12.12 0.04 1.09 (1.52) 12.18 0.31
0.05 1.25 (5.66) 2.46 1.76 -0.09 (-0.12) 15.59 -0.02 1.34 (1.86) 11.90 0.39
0.1 1.28 (5.73) 2.48 1.79 -0.26 (-0.35) 12.87 -0.07 1.53 (2.13) 11.78 0.45
0.2 1.34 (6.04) 2.48 1.87 -0.57 (-0.78) 12.34 -0.16 1.91 (2.68) 11.61 0.57
0.3 1.46 (6.54) 2.58 1.96 -0.91 (-1.27) 11.68 -0.27 2.37 (3.36) 11.09 0.74
0.4 1.88 (7.44) 3.16 2.06 -1.34 (-1.91) 11.05 -0.42 3.22 (4.89) 9.96 1.12
0.5 2.88 (6.97) 5.60 1.78 -1.82 (-2.68) 10.17 -0.62 4.70 (8.25) 7.17 2.27
0.6 3.44 (5.68) 8.70 1.37 -2.12 (-3.51) 8.35 -0.88 5.56 (9.41) 4.86 3.96
0.7 3.48 (4.73) 11.16 1.08 -1.42 (-3.57) 4.47 -1.10 4.90 (6.57) 8.53 1.99
0.8 2.98 (3.72) 12.59 0.82 0.44 (1.65) 2.46 0.62 2.55 (3.16) 11.78 0.75
0.9 2.12 (2.57) 13.35 0.55 0.75 (3.29) 2.36 1.10 1.37 (1.69) 12.49 0.38
0.95 1.71 (2.11) 13.16 0.45 0.86 (4.04) 2.44 1.22 0.85 (1.07) 12.80 0.23
0.99 1.36 (1.70) 13.09 0.36 0.97 (5.04) 2.75 1.22 0.39 (0.50) 12.28 0.11
Liquid Sample
Long Short Long-short
$\tau$ Mean Vol SR Mean Vol SR Mean Vol SR
0.01 0.95 (5.26) 2.99 1.10 0.16 (0.21) 11.09 0.05 0.79 (1.06) 11.40 0.24
0.05 0.96 (5.63) 2.89 1.15 0.15 (0.20) 12.99 0.04 0.81 (1.11) 11.22 0.25
0.1 0.98 (5.72) 2.90 1.17 0.11 (0.14) 12.70 0.03 0.87 (1.20) 11.59 0.26
0.2 1.02 (5.89) 2.92 1.21 0.04 (0.06) 13.86 0.01 0.98 (1.34) 11.32 0.30
0.3 1.10 (6.09) 3.05 1.25 -0.04 (-0.06) 13.86 -0.01 1.14 (1.58) 10.97 0.36
0.4 1.39 (6.31) 3.59 1.34 -0.14 (-0.20) 12.12 -0.04 1.53 (2.23) 9.81 0.54
0.5 1.86 (5.61) 5.86 1.10 -0.35 (-0.53) 10.10 -0.12 2.21 (4.21) 6.33 1.21
0.6 1.74 (3.69) 8.49 0.71 -0.38 (-0.88) 6.27 -0.21 2.11 (5.40) 5.45 1.34
0.7 1.66 (2.86) 10.27 0.56 0.48 (2.50) 3.08 0.54 1.18 (2.23) 9.51 0.43
0.8 1.28 (1.94) 11.37 0.39 0.71 (3.95) 2.89 0.85 0.57 (0.91) 10.97 0.18
0.9 0.91 (1.24) 11.68 0.27 0.76 (4.29) 2.83 0.93 0.15 (0.21) 10.39 0.05
0.95 0.77 (1.03) 11.60 0.23 0.80 (4.62) 2.83 0.98 -0.03 (-0.04) 10.39 -0.01
0.99 0.60 (0.80) 12.23 0.17 0.84 (4.72) 2.88 1.01 -0.23 (-0.32) 11.38 -0.07
@@
The average monthly returns and Sharpe ratios are the highest for the long-short portfolios formed based on the central $\tau$-quantiles, $\tau$s of 505050% or 606060%, for both the full sample and the liquid sample. Average monthly returns decrease as we move towards the tails of the distribution. For the full sample, average monthly returns are not significantly different from zero for $\tau$s greater than 80% and less than 10%. For the liquid sample, the average monthly returns fall even more sharply as one moves towards the tails of the distribution, becoming not significantly different from zero for $\tau$s greater than 70% and less than 40%. The highest Sharpe ratios are achieved by the long-short portfolios with the highest average monthly returns.
@@ 3.4.2 Distributional moments and the cross-section of stock returns
Table 8: Mean and median portfolios. This table shows average monthly returns and annualised Sharpe ratios for decile long-short equal-weighted portfolios based on the median and mean predictions from the two-stage quantile NN model, the mean prediction from the two-hidden-layer neural network with multi-head quantile loss (2hNN Quantile - Mean) and the mean prediction from the two-hidden-layer neural network with MSE loss (2hNN MSE - Mean). The results are shown for the full and liquid samples and for individual regions, covering the period from 1995 to 2018. The long portfolio is formed by taking the top 10% of stocks with the highest forecast each month, and the short portfolio is formed by taking the bottom 10% of stocks with the lowest forecast each month. Values in parentheses are t-statistics adjusted for Newey-West standard errors with 12 lags. The significance of the difference in mean returns is tested using t-statistics with Newey-West standard errors with 12 lags.
@@
To benchmark our forecasts, we use the mean forecasts obtained directly from the feed-forward neural network model estimated with the mean squared error loss function, as well as its distributional forecast equivalent where the mean forecast is derived from predicted quantiles. The long-short portfolio based on the mean forecast of our two-stage quantile model leads to significantly higher average monthly returns and Sharpe ratios compared to the portfolios based on the mean forecasts obtained directly from the simple neural network for the full sample across all regions. This is consistent with the $R^{2}$ results in Table 5, as our two-stage model outperforms the direct mean forecast in terms of $R^{2}$ across all regions and also leads to higher corresponding profitability. In the case of the full sample, the average monthly returns of our mean forecast are economically higher than the direct mean forecast in all regions and significantly higher for Japan, Asia-Pacific and globally.
Next, the average monthly returns of the mean- and median-based long-short portfolios are statistically indistinguishable for the liquid sample. This is also consistent with the $R^{2}$ results in Table 5, as there are no significant differences in the quality of the mean and median forecasts for the liquid sample. However, the results are different for the full sample. The difference between the mean and the median is associated with distributional asymmetry, which is more pronounced for stocks in the full sample. Mean-based long-short portfolios have significantly higher average monthly returns than median-based long-short portfolios. This is true for all regions and also globally. The difference between mean and median pricing for the full sample indicates the effect of forecast distributional asymmetry in the cross-section of stock returns.
@@
The relationship between predicted volatility and future returns is negative in the liquid sample but statistically insignificant, reflecting the overall ambiguity documented in prior studies on the pricing of volatility risk (Ang et al., 2006; Bollerslev et al., 2020). This result is consistent with the focus on large-cap stocks, where volatility effects are typically weaker. In the full sample, microstructure-related distortions in equal-weighted portfolios further obscure the relationship, although the pattern becomes somewhat clearer with value weighting.
@@ 4 Conclusion
We develop a novel machine learning approach to forecasting the distribution of stock returns, focusing on both US and international markets. Our method uses a multi-head quantile neural network to predict a set of $\tau$ quantiles based on stock characteristics and market variables. We interpolate these quantiles using cubic B-splines to derive the distribution and its moments, such as mean, volatility, skewness and kurtosis. This approach outperforms traditional parametric models and simpler neural networks in forecasting stock returns, providing more accurate out-of-sample mean and variance forecasts.
@@ B.1 Feed-forward neural networks
We use multi-head feed-forward neural networks with varying architectures as benchmarks for the two-stage model quantile forecasts. We consider a simple multi-output linear model without any hidden layers, neural networks with 1 hidden layer, and neural networks with 2 hidden layers as benchmarks for the two-stage model. Each model has 176 features and 37 outputs. The input features are the same as those for the standardised $\tau$-quantiles network in the two-stage model. The 37 outputs are the $\tau$-quantile forecasts of raw future stock returns for the same $\tau$ values as in the two-stage model. See Section 3.1.1 for the feature specification and Section 2 for the main model introduction.
[…]
@@
(b) One hidden layer neural network (1hNN)
@@
(c) Two hidden layers neural network (2hNN)
@@
(d) Single output neural network with two hidden layers (2hNN MSE)
@@
1(a) depicts the linear model (LNN), 1(b) illustrates the neural network with one hidden layer (1hNN), 1(c) shows the neural network with two hidden layers (2hNN), and 1(d) presents the two hidden layers neural network with a single output node (2hNN MSE), which is used to generate mean forecasts directly using the mean squared error loss function. All networks have 176 input features; multi-output networks have 37 outputs, varying in the number of hidden layers and neurons.
@@ B.2 GARCH benchmark
We consider several variants of GARCH(1,1) model with the standardised residuals distributed according to normal and t-distributions. Mean $\mu$ is either estimated or set equal to the risk-free rate at the end of the estimation window plus 5% divided by 252. Assuming these distributions, we simulate 100,000 paths of residuals for 22 days ahead and consequently derive the corresponding return paths. During this process, volatility is simulated using Equation 8
@@
For the t-distribution, besides estimating the number of degrees of freedom from the data, we also consider fixing the number of degrees of freedom to 3, 4, and 5. Furthermore, we try GARCH(2, 2) and EGARCH(1, 1, 1) specifications also with t-distributed standardised residuals, in both cases estimating degrees of freedom from the data.